assume the following economy

autonomous consumption =£3m

marginal propensity to consume= 0.8

business investment = 52

government spending =£127m

income tax rate = 40%


A. Find equilibrium size of income Y.

B. Identify whether the government runs a budget deficit or a budget surplus



1
Expert's answer
2021-04-16T07:28:25-0400

A. 

Equilibrium income 

"Y = C + I + G"

"Y = Ca + cYd + I + G"   

where "Yd = Y - tY"

"Y = \u00a33m + 0.8(Y - 0.4Y) + \u00a352m + \u00a3127m"

"Y- 0.48Y = \u00a3182m"  

"0.52Y = \u00a3182m"  

"Y =\frac{ £182m}{0.52}"  

"Y = \u00a3350m"  


B. Budget surplus "= tY - G"  

  "= 0.4(\u00a3350m) - \u00a3127m"

  "= \u00a3140m - \u00a3127m"

  "= \u00a313m."